# Hierarchical Retrieval with Evidence Curation for Open-Domain Financial Question Answering on Standardized Documents

### v2

@@ 1 Introduction
Retrieval-augmented generation (RAG)  Lewis et al. (2020) with large language models (LLMs) have significantly improved performance in knowledge-intensive tasks. Due to its ability to improve both factual accuracy and timeliness, extensive research  Yepes et al. (2024); Sarmah et al. (2024) has investigated applying RAG in searching financial information where accurate and up-to-date information is crucial for decision-making.
@@
To evaluate our approach, we assess QA performance in an open-domain setting. Existing financial question-answering benchmarks Islam et al. (2023); Lai et al. (2024) rely on small-scale corpora that include at most about 1,300 documents and very limited test sets, which do not reflect realistic financial scenarios. We propose LOFin (Large-scale Open-domain Financial QA), a comprehensive financial question-answering benchmark built on a large-scale corpus containing approximately 145,000 SEC filings from companies in the S&P 500. LOFin includes 1,595 open-domain question-answering test instances and addresses challenges in standardized document retrieval, such as near-duplicate tables and repetitive narratives, that are not evident in smaller datasets. In addition, we release the entire benchmark as open-source to support future research in the field.
Experimental results indicate that HiREC improves performance by at least 13% compared to existing RAG methods. In addition, our framework consistently outperforms commercial llms with web search engines such as SearchGPT OpenAI (2025) and Perplexity Perplexity (2023).
@@ 2 Related Works
Financial QA has advanced significantly, with recent benchmarks emphasizing numeric reasoning and table understanding. TAT-QA Zhu et al. (2021) and FinQA Chen et al. (2021) provide single-page tabular contexts, while DocFinQA Reddy et al. (2024) and DocMath-Eval Zhao et al. (2024) extend this to multi-page settings. However, these benchmarks remain closed-domain, limiting their applicability for RAG systems. Open-domain benchmarks have also been proposed. For instance, consider Financebench Islam et al. (2023) and SEC-QA Lai et al. (2024). However, these datasets are built on small-scale document collections and suffer from limited test set sizes or the lack of publicly available fixed test sets.
With the emergence of financial QA, research on financial RAG has also been progressing. There are graph-based methods Barry et al. (2025); Sarmah et al. (2024) tailored to the financial domain as well as hybrid approaches Wang (2024). In addition, studies on financial document chunking offer valuable insights Yepes et al. (2024).
@@
RAG has evolved in diverse ways Gao et al. (2023); Zhang et al. (2025). Hierarchical retrieval was proposed for cases where sections are clearly demarcated, typically employing a two-step document-passage process Arivazhagan et al. (2023); Chen et al. (2024). The difference from previous studies is that while they segment documents into sections, we segment at the level of individual documents. There are also studies that utilize filtering to enhance the quality of the contexts input into RAG Zhuang et al. (2024); Wang et al. (2024b). Unlike previous approaches that required training, we manage quality using only an LLM. Iterative retrieval is typically proposed for multi-hop QA. A standard iterative method uses the context retrieved in the first step as part of the query for subsequent iterations Trivedi et al. (2022); Shao et al. (2023). Self-RAG Asai et al. (2023) also conducts iterative retrieval that includes the generation process. However, our approach does not utilize previously retrieved context because we specifically need to discover missing information.
@@ 3.1 Large-scale Document Collection
We converted the HTML documents to PDF using the wkhtmltopdf [^2] https://wkhtmltopdf.org/ library. Following the approach of Islam et al., 2023, we used the PyMuPDF library [^3] https://pymupdf.readthedocs.io/ to extract text at the page level from these PDFs, excluding reports that lacked proper page separation. The final corpus consists of 145,897 reports from 516 companies.
@@ 3.2 Open-domain QA Pair Construction
In this section, we detail our process for constructing open-domain QA pairs by leveraging three established financial QA benchmarks: FinQA Chen et al. (2021), Financebench Islam et al. (2023), and SEC-QA Lai et al. (2024).
@@ 4.1.1 Document Retriever
1:  $\mathcal{P}_{f}\leftarrow\emptyset$ // Initialization
2:  $\mathcal{P}_{r}\leftarrow$ HierarchicalRetrieval$(q,\mathcal{D})$eqparboxCOMMENT// Sec 4.1
@@
11:  return  AnswerGeneration$(q,\mathcal{P}_{r})$eqparboxCOMMENT// Sec 4.3
Documents contain extensive context, and their standardized format makes it difficult to capture all important details in a single vector. To address this, we extract and index key distinguishing information when retrieving documents. In financial reports, the cover page provides essential details such as the company name, report type, and fiscal period. For each document $d\in\mathcal{D}$, we generate a cover page summary $d^{\prime}$ using an LLM (the prompt is detailed in Appendix D.1), precompute its embedding with a bi-encoder Wang et al. (2024a), and index the resulting vector as $\mathbf{v}_{d}=E^{D}(d^{\prime})$ in the document store.
@@
Our document retrieval process consists of three stages. First, given a question $q$, we use an LLM to convert it into a refined query $q^{\prime}$. This conversion reduces issues caused by extraneous financial terms (e.g., tickers or service names like Google or Facebook) that may hinder effective retrieval. (Details of the transformation prompt are provided in Appendix D.2.) Second, we perform dense retrieval. We compute the vector representation of $q^{\prime}$ using the same bi-encoder employed during document indexing, yielding $\mathbf{v}_{q^{\prime}}=E^{D}(q^{\prime})$. The relevance between $q^{\prime}$ and each document $d$ is then determined via $s^{D}_{q^{\prime},d}=\mathbf{v}_{q^{\prime}}^{\top}\mathbf{v}_{d}$ Karpukhin et al. (2020); based on these scores, we retrieve $k^{\prime}_{D}$ candidate documents $\mathcal{D}_{\text{cand}}$. Finally, we rerank these candidates using a cross-encoder He et al. (2021) by computing $\text{CrossEncoder}^{D}(q^{\prime},d)$ and select the top $k_{D}$ documents $\mathcal{D}_{r}$. This multi-stage process ultimately produces the final set of documents $\mathcal{D}_{r}$ that are most relevant to the original question $q$.
@@ 4.1.2 Passage Retriever
Within the final set of retrieved documents $\mathcal{D}_{r}$, the passage retriever evaluates each passage $p$ by computing a score using a cross-encoder, $\text{CrossEncoder}^{P}(q,p)$. It then selects the top $k_{P}$ passages to form the final set $\mathcal{P}_{r}$. By reducing the number of passages processed by the cross-encoder, we enable real-time computation while still taking advantage of its superior ability to capture inter-sentence relationships compared to a bi-encoder. However, passage retrievers that are pretrained on general text typically have difficulty handling financial tables. For instance, when retrieving tables, attributes such as titles, periods, and indicators are more important than the numerical values, yet these cues are not well captured. To address this, we fine-tuned the model on table data using FinQA training set where tables serve as evidence. Specifically, for each question $q$ and its associated evidence document $d$, we denote the tables on the evidence page as the evidence passage set $\mathcal{X}$. For each evidence passage $p\in\mathcal{X}$, we sample $n_{neg}$ negative passages, where negative passages are defined as tables that appear on pages other than the evidence page. The cross-encoder is then trained with a binary cross-entropy loss Nogueira and Cho (2019) defined as
@@ 4.2 Evidence Curation
Financial questions often involve comparisons between different time periods or companies. Even when the retrieval process selects relevant passages, some critical information may be missing. Furthermore, retrieved passages can contain irrelevant data that hinders overall performance Liu et al. (2024); Xu et al. (2024). To overcome these issues, we introduce an evidence curation process that filters out incorrect data and fills information gaps by initiating additional retrieval when necessary. Our process comprises three modules: a passage filter that removes irrelevant passages (Section 4.2.1), an answerability checker that assesses evidence sufficiency (Section 4.2.2), and a complementary question generator that formulates a supplementary query if necessary (Section 4.2.3). We use an LLM to perform all three tasks in a single response (see Appendix D.3 for the prompt).
@@ 4.3 Answer Generation
In the answer generation stage, the relevant passages and the original question serve as inputs to a reasoning process that derives the final answer. For questions requiring numerical calculations, a Program-of-Thought (PoT) Chen et al. (2022) reasoning method is employed, while for text-based inferential questions, a Chain-of-Thought (CoT) Wei et al. (2022) approach is applied. This dual strategy is particularly effective for financial documents, which are rich in numerical data and tables, ensuring comprehensive and accurate reasoning. See Appendix D.4 for prompt details, adapted from DocMath-Eval Zhao et al. (2024).
@@
Method Page Recall Answer Acc Page Recall Answer Acc Page Recall Answer Acc Page Recall Answer Acc $\mathbf{k}$
@@
Perplexity Perplexity (2023) - 2.51 - 5.13 - 24.00 - 10.55 -
Self-RAG Asai et al. (2023) 19.96 1.86 24.18 4.03 12.75 17.00 18.96 7.63 10.0
RQ-RAG Chan et al. (2024) 18.61 1.97 19.05 2.56 17.96 20.50 18.54 8.34 36.0
IRCoT Trivedi et al. (2022) $\lozenge$ 28.17 19.10 34.62 27.84 12.67 20.00 25.15 22.31 20.0
HybridSearch Wang (2024) $\lozenge$ 26.75 19.10 32.05 30.77 14.37 27.00 24.39 25.62 10.0
HHR Arivazhagan et al. (2023) $\lozenge$ 37.67 26.53 40.29 32.97 21.98 26.50 33.31 28.67 10.0
Dense Karpukhin et al. (2020) $\lozenge$ 37.69 23.69 40.48 32.97 26.18 31.00 34.78 29.22 10.0
@@ 5.1 Experimental Settings
Our system leverages several pre-trained and fine-tuned models across its components. For the passage retriever, we fine-tune a DeBERTa-v3 model He et al. (2021) with $n_{neg}=8$ for negative sampling, a batch size of 128, 3 epochs, and a learning rate of $2\times 10^{-7}$ on a single GeForce RTX 4090 GPU. For document retrieval, we employ the E5 model Wang et al. (2024a) as the bi-encoder and use DeBERTa-v3 as the reranker. The answer generator is powered by OpenAI’s GPT-4o, while other LLM-based tasks (query transformation, document summarization, and evidence curation) are handled by Qwen-2.5-7B-Instruct Yang et al. (2024).
@@
We compare our proposed method with several state-of-the-art RAG approaches, including RQ-RAG Chan et al. (2024) and Self-RAG Asai et al. (2023). To ensure a fair comparison with our approach—which employs GPT-4o as the answer generator—we use the same answer generator for the latest retrieval algorithms: Dense, HybridSearch Wang (2024), HHR Arivazhagan et al. (2023), and IRCoT Trivedi et al. (2022). In addition, we evaluate against Perplexity Perplexity (2023), a commercial LLM service that combines web search with LLMs. RQ-RAG, Self-RAG, and IRCoT employ iterative LLM-based retrieval, while Dense serves as a strong baseline, it employs OpenAI’s text-embedding-3-small as its encoder for DPR, with results reranked by DeBERTa-v3. All methods construct passages by concatenating the title and content during retrieval. Detailed configurations for each baseline are provided in Appendix C.2.
@@
We evaluate numeric answers for accuracy by considering rounding and truncation, while textual answers are evaluated using GPT-4o and FAMMA prompts Xue et al. (2024) (see Appendix C.3). We measure retrieval performance at the page level, using recall and precision against ground-truth evidence pages as metrics. In particular, since chunk locations or units can vary for Page recall, we standardize them at the page level to ensure consistent performance measurement.
@@ 5.2 Main Result
Method Page Precision Page Recall Ansswer Accuracy
@@ 5.3 Analysis
HR is critical for enhancing retrieval accuracy because its absence produces the lowest performance. In addition, the initial search performance of HR exceeds that of the dense method combined with EC. The w/o EC setting demonstrates the effectiveness of both passage filter and the complementary question generator. HR supported by EC delivers enhanced precision, recall and answer accuracy. Furthermore, results from the w/o Filter setting reveal the advantages of the complement and filtering. Although the complementary component attains the highest recall score in the absence of filtering, its accuracy is lower than that of HiREC with filtering. This outcome is attributed to the inclusion of incorrect information that causes conflicts when filtering is not applied Xu et al. (2024).
@@
Dense - - 2,666 / 177 9.3 / 2.5
IRCoT 9,610 / 1,372 33.4 / 19.1 3,475 / 177 12.1 / 2.5
HiREC 4,291 / 313 14.9 / 4.4 1,052 / 159 3.7 / 2.2
@@ 5.4 Analysis of Performance Across Various LLM Generators
Method Numeric (Table) Numeric (Text) Textual Average
@@
We thoroughly analyze the effectiveness of the HiREC framework across diverse LLM generator models. For this analysis, additional evaluations were conducted using open-source models such as DeepSeek-R1-Distill-Qwen-14BGuo et al. (2025) and Qwen-2.5-7B-InstructYang et al. (2024) as generators, in place of GPT-4o.
@@
Page Recall Answer Acc Page Recall Answer Acc Page Recall Answer Acc
@@
Evidence
Curation Relevant passages: ADBE_2015_10K (p.59)  Answerable: False
@@
Evidence
Curation Relevant passages: ADBE_2015_10K (p.59), ADBE_2016_10K (p.61)  Answerable: True

### v3

@@ 1 Introduction
Retrieval-augmented generation (RAG)  lewis2020retrieval with large language models (LLMs) have significantly improved performance in knowledge-intensive tasks. Due to its ability to improve both factual accuracy and timeliness, extensive research  yepes2024financial; sarmah2024hybridrag has investigated applying RAG in searching financial information where accurate and up-to-date information is crucial for decision-making.
@@
To evaluate our approach, we assess QA performance in an open-domain setting. Existing financial question-answering benchmarks islam2023financebench; lai2024sec rely on small-scale corpora that include at most about 1,300 documents and very limited test sets, which do not reflect realistic financial scenarios. We propose LOFin (Large-scale Open-domain Financial QA), a comprehensive financial question-answering benchmark built on a large-scale corpus containing approximately 145,000 SEC filings from companies in the S&P 500. LOFin includes 1,595 open-domain question-answering test instances and addresses challenges in standardized document retrieval, such as near-duplicate tables and repetitive narratives, that are not evident in smaller datasets. In addition, we release the entire benchmark as open-source to support future research in the field.
Experimental results indicate that HiREC improves performance by at least 13% compared to existing RAG methods. In addition, our framework consistently outperforms commercial llms with web search engines such as SearchGPT searchGPT and Perplexity Perplexity2023.
@@ 2 Related Works
Financial QA has advanced significantly, with recent benchmarks emphasizing numeric reasoning and table understanding. TAT-QA zhu2021tat and FinQA chen2021finqa provide single-page tabular contexts, while DocFinQA reddy2024docfinqa and DocMath-Eval zhao-etal-2024-docmath extend this to multi-page settings. However, these benchmarks remain closed-domain, limiting their applicability for RAG systems. Open-domain benchmarks have also been proposed. For instance, consider Financebench islam2023financebench and SEC-QA lai2024sec. However, these datasets are built on small-scale document collections and suffer from limited test set sizes or the lack of publicly available fixed test sets.
With the emergence of financial QA, research on financial RAG has also been progressing. There are graph-based methods barry2025graphrag; sarmah2024hybridrag tailored to the financial domain as well as hybrid approaches wang2024financerag. In addition, studies on financial document chunking offer valuable insights yepes2024financial.
@@
RAG has evolved in diverse ways gao2023retrieval; zhang2025survey. Hierarchical retrieval was proposed for cases where sections are clearly demarcated, typically employing a two-step document-passage process arivazhagan-etal-2023-hybrid; chen2024hiqa. The difference from previous studies is that while they segment documents into sections, we segment at the level of individual documents. There are also studies that utilize filtering to enhance the quality of the contexts input into RAG zhuang-etal-2024-efficientrag; wang2024rear. Unlike previous approaches that required training, we manage quality using only an LLM. Iterative retrieval is typically proposed for multi-hop QA. A standard iterative method uses the context retrieved in the first step as part of the query for subsequent iterations trivedi2022interleaving; shao2023enhancing. Self-RAG asai2023self also conducts iterative retrieval that includes the generation process. However, our approach does not utilize previously retrieved context because we specifically need to discover missing information.
@@ 3.1 Large-scale Document Collection
We converted the HTML documents to PDF using the wkhtmltopdf [^2] https://wkhtmltopdf.org/ library. Following the approach of islam2023financebench, we used the PyMuPDF library [^3] https://pymupdf.readthedocs.io/ to extract text at the page level from these PDFs, excluding reports that lacked proper page separation. The final corpus consists of 145,897 reports from 516 companies.
@@ 3.2 Open-domain QA Pair Construction
In this section, we detail our process for constructing open-domain QA pairs by leveraging three established financial QA benchmarks: FinQA chen2021finqa, Financebench islam2023financebench, and SEC-QA lai2024sec.
@@ 4.1.1 Document Retriever
[…]
@@
[…]
Documents contain extensive context, and their standardized format makes it difficult to capture all important details in a single vector. To address this, we extract and index key distinguishing information when retrieving documents. In financial reports, the cover page provides essential details such as the company name, report type, and fiscal period. For each document $d\in\mathcal{D}$, we generate a cover page summary $d^{\prime}$ using an LLM (the prompt is detailed in Appendix D.1), precompute its embedding with a bi-encoder wang2024multilingual, and index the resulting vector as $\mathbf{v}_{d}=E^{D}(d^{\prime})$ in the document store.
@@
Our document retrieval process consists of three stages. First, given a question $q$, we use an LLM to convert it into a refined query $q^{\prime}$. This conversion reduces issues caused by extraneous financial terms (e.g., tickers or service names like Google or Facebook) that may hinder effective retrieval. (Details of the transformation prompt are provided in Appendix D.2.) Second, we perform dense retrieval. We compute the vector representation of $q^{\prime}$ using the same bi-encoder employed during document indexing, yielding $\mathbf{v}_{q^{\prime}}=E^{D}(q^{\prime})$. The relevance between $q^{\prime}$ and each document $d$ is then determined via $s^{D}_{q^{\prime},d}=\mathbf{v}_{q^{\prime}}^{\top}\mathbf{v}_{d}$ karpukhin2020dense; based on these scores, we retrieve $k^{\prime}_{D}$ candidate documents $\mathcal{D}_{\text{cand}}$. Finally, we rerank these candidates using a cross-encoder he2021debertav3 by computing $\text{CrossEncoder}^{D}(q^{\prime},d)$ and select the top $k_{D}$ documents $\mathcal{D}_{r}$. This multi-stage process ultimately produces the final set of documents $\mathcal{D}_{r}$ that are most relevant to the original question $q$.
@@ 4.1.2 Passage Retriever
Within the final set of retrieved documents $\mathcal{D}_{r}$, the passage retriever evaluates each passage $p$ by computing a score using a cross-encoder, $\text{CrossEncoder}^{P}(q,p)$. It then selects the top $k_{P}$ passages to form the final set $\mathcal{P}_{r}$. By reducing the number of passages processed by the cross-encoder, we enable real-time computation while still taking advantage of its superior ability to capture inter-sentence relationships compared to a bi-encoder. However, passage retrievers that are pretrained on general text typically have difficulty handling financial tables. For instance, when retrieving tables, attributes such as titles, periods, and indicators are more important than the numerical values, yet these cues are not well captured. To address this, we fine-tuned the model on table data using FinQA training set where tables serve as evidence. Specifically, for each question $q$ and its associated evidence document $d$, we denote the tables on the evidence page as the evidence passage set $\mathcal{X}$. For each evidence passage $p\in\mathcal{X}$, we sample $n_{neg}$ negative passages, where negative passages are defined as tables that appear on pages other than the evidence page. The cross-encoder is then trained with a binary cross-entropy loss nogueira2019passage defined as
@@ 4.2 Evidence Curation
Financial questions often involve comparisons between different time periods or companies. Even when the retrieval process selects relevant passages, some critical information may be missing. Furthermore, retrieved passages can contain irrelevant data that hinders overall performance liu2024lost; xu2024knowledge. To overcome these issues, we introduce an evidence curation process that filters out incorrect data and fills information gaps by initiating additional retrieval when necessary. Our process comprises three modules: a passage filter that removes irrelevant passages (Section 4.2.1), an answerability checker that assesses evidence sufficiency (Section 4.2.2), and a complementary question generator that formulates a supplementary query if necessary (Section 4.2.3). We use an LLM to perform all three tasks in a single response (see Appendix D.3 for the prompt).
@@ 4.3 Answer Generation
In the answer generation stage, the relevant passages and the original question serve as inputs to a reasoning process that derives the final answer. For questions requiring numerical calculations, a Program-of-Thought (PoT) chen2022program reasoning method is employed, while for text-based inferential questions, a Chain-of-Thought (CoT) wei2022chain approach is applied. This dual strategy is particularly effective for financial documents, which are rich in numerical data and tables, ensuring comprehensive and accurate reasoning. See Appendix D.4 for prompt details, adapted from DocMath-Eval zhao-etal-2024-docmath.
@@
Method
Page
Recall
Answer
Acc
Page
Recall
Answer
Acc
Page
Recall
Answer
Acc
Page
Recall
Answer
Acc
$\mathbf{k}$
@@
Perplexity Perplexity2023 - 2.51 - 5.13 - 24.00 - 10.55 -
Self-RAG asai2023self 19.96 1.86 24.18 4.03 12.75 17.00 18.96 7.63 10.0
RQ-RAG chan2024rq 18.61 1.97 19.05 2.56 17.96 20.50 18.54 8.34 36.0
IRCoT trivedi2022interleaving $\lozenge$ 28.17 19.10 34.62 27.84 12.67 20.00 25.15 22.31 20.0
HybridSearch wang2024financerag $\lozenge$ 26.75 19.10 32.05 30.77 14.37 27.00 24.39 25.62 10.0
HHR arivazhagan-etal-2023-hybrid $\lozenge$ 37.67 26.53 40.29 32.97 21.98 26.50 33.31 28.67 10.0
Dense karpukhin2020dense $\lozenge$ 37.69 23.69 40.48 32.97 26.18 31.00 34.78 29.22 10.0
@@ 5.1 Experimental Settings
Our system leverages several pre-trained and fine-tuned models across its components. For the passage retriever, we fine-tune a DeBERTa-v3 model he2021debertav3 with $n_{neg}=8$ for negative sampling, a batch size of 128, 3 epochs, and a learning rate of $2\times 10^{-7}$ on a single GeForce RTX 4090 GPU. For document retrieval, we employ the E5 model wang2024multilingual as the bi-encoder and use DeBERTa-v3 as the reranker. The answer generator is powered by OpenAI’s GPT-4o, while other LLM-based tasks (query transformation, document summarization, and evidence curation) are handled by Qwen-2.5-7B-Instruct yang2024qwen2.
@@
We compare our proposed method with several state-of-the-art RAG approaches, including RQ-RAG chan2024rq and Self-RAG asai2023self. To ensure a fair comparison with our approach—which employs GPT-4o as the answer generator—we use the same answer generator for the latest retrieval algorithms: Dense, HybridSearch wang2024financerag, HHR arivazhagan-etal-2023-hybrid, and IRCoT trivedi2022interleaving. In addition, we evaluate against Perplexity Perplexity2023, a commercial LLM service that combines web search with LLMs. RQ-RAG, Self-RAG, and IRCoT employ iterative LLM-based retrieval, while Dense serves as a strong baseline, it employs OpenAI’s text-embedding-3-small as its encoder for DPR, with results reranked by DeBERTa-v3. All methods construct passages by concatenating the title and content during retrieval. Detailed configurations for each baseline are provided in Appendix C.2.
@@
We evaluate numeric answers for accuracy by considering rounding and truncation, while textual answers are evaluated using GPT-4o and FAMMA prompts xue2024famma (see Appendix C.3). We measure retrieval performance at the page level, using recall and precision against ground-truth evidence pages as metrics. In particular, since chunk locations or units can vary for Page recall, we standardize them at the page level to ensure consistent performance measurement.
@@ 5.2 Main Result
Method
Page
Precision
Page
Recall
Ansswer
Accuracy
@@ 5.3 Analysis
HR is critical for enhancing retrieval accuracy because its absence produces the lowest performance. In addition, the initial search performance of HR exceeds that of the dense method combined with EC. The w/o EC setting demonstrates the effectiveness of both passage filter and the complementary question generator. HR supported by EC delivers enhanced precision, recall and answer accuracy. Furthermore, results from the w/o Filter setting reveal the advantages of the complement and filtering. Although the complementary component attains the highest recall score in the absence of filtering, its accuracy is lower than that of HiREC with filtering. This outcome is attributed to the inclusion of incorrect information that causes conflicts when filtering is not applied xu2024knowledge.
@@
Dense - -
2,666
/ 177
9.3
/ 2.5
IRCoT
9,610
/ 1,372
33.4
/ 19.1
3,475
/ 177
12.1
/ 2.5
HiREC
4,291
/ 313
14.9
/ 4.4
1,052
/ 159
3.7
/ 2.2
@@ 5.4 Analysis of Performance Across Various LLM Generators
Method
Numeric
Numeric
Textual Average
@@
We thoroughly analyze the effectiveness of the HiREC framework across diverse LLM generator models. For this analysis, additional evaluations were conducted using open-source models such as DeepSeek-R1-Distill-Qwen-14Bguo2025deepseek and Qwen-2.5-7B-Instructyang2024qwen2 as generators, in place of GPT-4o.
@@
Page
Recall
Answer
Acc
Page
Recall
Answer
Acc
Page
Recall
Answer
Acc
@@
Evidence Curation Relevant passages: ADBE_2015_10K (p.59) Answerable: False
@@
Evidence Curation Relevant passages: ADBE_2015_10K (p.59), ADBE_2016_10K (p.61)  Answerable: True